OpenStudy (anonymous):

how much money must be deposited in an account paying 7.25% annual interest, compounded quarterly, to have a balance of $1000 after 10 years?

OpenStudy (dumbcow):

General formula: B = D(1+i/n)^nt i = int rate n = num of times per year interest credited t=num years

OpenStudy (anonymous):

what do you mean b=d?

OpenStudy (dumbcow):

b = ending balance d = initial deposit

OpenStudy (anonymous):

im confused

OpenStudy (dumbcow):

ok lets say you deposit $100 and earn 50% 50% of 100 is 50 so your balance at end of year is 150 so 150=100*(1+0.5)

OpenStudy (anonymous):

Compound Formula for calculating compound interest: Where, \[A=P(1+r/t)^{nt}\] A = final amount P = principal amount (initial investment) r = annual nominal interest rate (as a decimal) n = number of times the interest is compounded per year t = number of years

OpenStudy (anonymous):

so just plug in the 1000, the interest rate .0725, number of years and compounds per year

OpenStudy (anonymous):

i do not know how i would set it up like to plug it in or anything

OpenStudy (dumbcow):

1000=D*(1+.0725/4)^40 solve for D

OpenStudy (anonymous):

\[1000/(1+.0725/4)^{40}=P\] That is what you end up with.

OpenStudy (anonymous):

solve for t or p?

OpenStudy (anonymous):

There is no t to solve for

OpenStudy (anonymous):

You already know t, its 10 years

OpenStudy (anonymous):

so solve for p?

OpenStudy (anonymous):

It's already solved, you just have to plug it into a calculator

OpenStudy (anonymous):

or do it by hand if you want to.

OpenStudy (anonymous):

1000/(40725/40000)^40 = 487.48